# DynaFlow: Dynamics-embedded Flow Matching for Physically Consistent Motion Generation from State-only Demonstrations

### v2

@@ Abstract
This paper introduces DynaFlow, a novel framework that embeds a differentiable simulator directly into a flow matching model. By generating trajectories in the action space and mapping them to dynamically feasible state trajectories via the simulator, DynaFlow ensures all outputs are physically consistent by construction. This end-to-end differentiable architecture enables training on state-only demonstrations, allowing the model to simultaneously generate physically consistent state trajectories while inferring the underlying action sequences required to produce them. We demonstrate the effectiveness of our approach through quantitative evaluations and showcase its real-world applicability by deploying the generated actions onto a physical Go1 quadruped robot. The robot successfully reproduces diverse gait present in the dataset, executes long-horizon motions in open-loop control and translates infeasible kinematic demonstrations into dynamically executable, stylistic behaviors. These hardware experiments validate that DynaFlow produces deployable, highly effective motions on real-world hardware from state-only demonstrations, effectively bridging the gap between kinematic data and real-world execution.
@@ I Introduction
However, directly applying these generative models to character animation and robot control presents significant challenges. A primary limitation is the lack of physical consistency. Most generative models learn a statistical approximation of the data distribution from a finite set of examples, rather than the underlying physical principles governing the data. Consequently, there is no guarantee that the generated outputs will adhere to intrinsic physical principles or dynamic constraints. This means that when generating novel behaviors, the resulting motions can be either physically implausible, with artifacts like ground penetration, character floating, and foot sliding, or dynamically inconsistent, making them unsuitable to execute in the physical world. The issue becomes particularly pronounced when models are trained on datasets with inherent physical inconsistencies, such as those sourced from motion capture or generated through kinematic retargeting.
Another major hurdle is the scarcity of action-labeled data. While state trajectories can often be obtained from motion capture or other sources, the corresponding action sequences (e.g., joint torques or motor commands) are rarely available and are costly to collect, typically being hardware-specific. Prior diffusion-based control approaches can be categorized into two principal ways. One option is to directly train a policy to predict actions from true action data, but this is rarely viable given the limited availability of true action data. Alternatively, a hierarchical framework can be employed, where the diffusion model generates desired state trajectories that are then executed by a low-level tracking controller. However, this hierarchical strategy has its own drawback: the tracking controller usually requires extensive fine-tuning for each new motion, and ensuring robustness against the distributional gap between planned trajectories and the controller remains a significant challenge [9].
@@
To overcome these limitations, we propose DynaFlow, a novel framework that guarantees physically consistent motion generation by embedding dynamics directly into the generation process, while simultaneously inferring actions from state-only demonstration data. The core idea of DynaFlow is to integrate a differentiable simulator at the output of the flow matching prediction module. This simulator layer acts as a mapping from the space of action trajectories to the space of dynamically feasible state trajectories, ensuring that the model’s output strictly adheres to the laws of physics by construction. Furthermore, its differentiable nature allows the entire model to be trained end-to-end. During this process, the model naturally discovers the action trajectory required to reconstruct a given state trajectory in a dynamically consistent manner, even without explicit action labels.
@@
We conduct a series of experiments to validate the effectiveness of DynaFlow. Our quantitative analysis compares DynaFlow against several baselines, evaluating both dynamic feasibility and distributional similarity on two distinct datasets: a rich, strictly feasible dataset of quadruped locomotion and a challenging single-trajectory dataset from retargeted motion capture. Our results demonstrate that DynaFlow consistently generates strictly feasible trajectories, even when trained on a physically inconsistent dataset, while remaining competitive in distributional similarity. To showcase its real-world applicability, we deploy action trajectories generated by DynaFlow on a physical Go1 quadruped robot. The robot successfully reproduces diverse gaits observed in the training data and executes long-horizon motions with high accuracy in challenging open-loop experiments, validating the precision and coherence of the generated actions. Furthermore, we demonstrate its ability to translate infeasible retargeted motions into dynamically executable and stylistic behaviors on hardware, bridging the gap between kinematic demonstration and real-world execution.
@@
  • Dynamics-embedded generative modeling: We propose DynaFlow, a novel generative model that embeds a differentiable simulator into a flow matching framework to guarantee all generated trajectories are strictly dynamically consistent by construction.
  • Learning from state-only data: Our method overcomes the common challenge of unavailable action data by leveraging analytical gradients, learning action directly from state-only demonstrations.
  • Real-world validation: We demonstrate the practical viability of our approach through successful hardware deployment of the generated motion on a physical quadruped robot, bridging the gap between motion generation and real-world execution.
@@ II-A Generative Models for Control Tasks
The first paradigm is the hierarchical approach, where the generative model serves as a high-level state planner, while a separate low-level control module produces executable actions. This factorized design decouples planning and control, offering the advantage of addressing the problem in two stages and enabling the generative model to focus on state trajectory generation while other control modules are used for execution. A key design choice in this paradigm is how to implement the control module. For example, Decision Diffuser [10] uses ground-truth action labels to train an inverse dynamics network via supervised learning. In the absence of action labels, one common approach is to pair the planner with a reinforcement learning-based tracking controller [11, 12]. Alternatively, model-based approaches combine generative models with optimization-based controllers, where diffusion models either generate intermediate subgoals that are tracked by Model Predictive Control (MPC)[13] or provide initial guesses that are refined by trajectory optimization[14]. While effective, these approaches are often limited by their reliance on an auxiliary low-level tracking module, which is often sensitive to out-of-distribution plans generated by the high-level planner[9] or requires extensive, task-specific tuning. This dependency hinders the scalability and generalization of hierarchical approaches.
The second paradigm is the direct action generation approach, where the generative model itself infers executable control inputs. Models in this category either model action trajectories directly or jointly model state–action pairs. This line of work has shown notable success across diverse domains, including manipulation [8, 15, 16], legged locomotion[17, 18], character control[9, 19, 20], and humanoid whole-body control[21]. While some works focus on generating action trajectories[8, 15, 16, 17, 20], others adopt joint formulations over state–action pairs [9, 21] or couple states with latent action representations [19]. Such formulations alleviate the challenge of action sequence conditioning and further demonstrate how task-level guidance can steer generated actions for downstream tasks. Despite these advances, the reliance on ground-truth action datasets poses a practical limitation, as such data is costly to obtain and often requires training or fine-tuning of expert controllers for each task[20].
@@ II-B Physics-Constrained Motion Generation
In contrast, our framework guarantees a physically consistent trajectory by construction, embedding the simulator within the generative process and removing the need for any external projection or guidance.
@@ II-C Differentiable Physics
The development of differentiable simulators[27, 28] has provided powerful tools for incorporating underlying physics directly into the learning process. By leveraging physically consistent gradients, these simulators enable sample-efficient training and provide interpretable learning signals from states. Building on this capability, several works have proposed policy learning methods that utilize gradient signals from differentiable simulators to guide policy optimization[29, 30]. Another line of research leverages the ability to backpropagate from states to actions, reformulating motion or skill imitation as a state-matching problem. DiffMimic[31] demonstrated efficient learning of dynamic character control from a single demonstration, while ILD[32] introduced a bidirectional state-matching objective that stabilizes training. In addition, differentiable simulation has been combined with state-diffusion planning to train latent action representations[33], which are then employed as tracking controllers for language-guided character control.
Our framework directly embeds a differentiable simulator into the flow matching framework, optimizing action trajectories to match a target probability flow in the state space. This tight integration enables end-to-end learning of physically-grounded generative policy directly from state-only demonstrations.
@@ IV Dynamics-embedded Flow Matching
Our proposed framework, DynaFlow, guarantees the dynamic feasibility of generated trajectory by embedding a simulator as a deterministic mapping module at the output, which projects preceding outputs into the space of physically valid state trajectories, ensuring feasibility by construction.
@@ IV-A Differentiable Physics as a Dynamics-Aware Embedding Layer
  • Action Prediction Network. A network $D_{\theta}$ predicts a sequence of future actions $U$ conditioned on the current noisy state trajectory $X_{t}$, auxiliary conditioning $c$ including observations, command signals, and the time step $t$:
@@ IV-B Training Objective
Here, the expectation is over the time $t\sim\mathcal{U}(0,1)$, the base distribution $X_{0}\sim p_{0}$, and the target data distribution represented by the random variable $X_{1}\sim p_{1}$. $X_{t}$ denotes their linear interpolation and $W$ is a weighting mask applied element-wise to balance contributions across state dimensions and timesteps. In practice, we approximate this expectation by sampling an expert trajectory $\tilde{X}_{1}\sim\mathcal{D}$ and generate a corresponding prediction $\hat{X}_{1}=\Phi(x_{0},D_{\theta}(X_{t},c,t))$.
@@ IV-D Datasets
We use two distinct datasets to evaluate performance under both controlled and challenging conditions. The first, Simulation Rollouts Dataset, comprises about 12,000 5-second trot and bound trajectories generated by rolling out a pre-trained policy [36] under various velocity commands. For quantitative evaluation, we additionally re-generated the Simulation Rollouts Dataset by rolling out the same pre-trained policy in MuJoCo[37] under identical condition, offering a rich set of dynamically feasible state–action data for controlled evaluation. The second, Retargeted Motion Capture Dataset, is a more challenging state-only dataset consisting of a single 2.54-second galloping trajectory. The motion was created by adapting German Shepherd motion capture data [38] to a Go1 robot skeleton, following the retargeting procedure of Peng et al. [39]. As kinematic retargeting does not incorporate dynamic properties or enforce actuation constraints, the resulting trajectory is often dynamically infeasible, exhibiting physical artifacts such as ground penetration. This makes it a demanding benchmark of model robustness and reflects real-world scenarios where only sparse kinematic data are available.
@@ IV-E Implementation Details
Our framework is implemented using JAX[40] combined with the massively parallel differentiable robotics simulation MuJoCo MJX[28].
@@
We use a 1D Diffusion Transformer (DiT) backbone with three transformer blocks (10.3M parameters). The model is conditioned with sinusoidal positional encodings over the horizon, MLP-based timestep embeddings, and task-specific attributes $c=[\,g,v,\omega,q,\dot{q},z_{\mathrm{gait}},v_{x}^{\mathrm{cmd}},v_{y}^{\mathrm{cmd}},\omega_{z}^{\mathrm{cmd}}]$. This vector includes gravity, base velocities, joint states, gait mode, and velocity commands. We train with Adam[41] optimizer with a learning rate of $2{\times}10^{-4}$, and apply Exponential Moving Averaging (EMA) with decay $0.995$ for 25k steps ($\approx$ 3h) on a single NVIDIA RTX 4090 GPU with a batch size of 1024.
@@ V-A Comparative Analysis on Physical Consistency and Distributional Similarity
Figure 3: Comparison of tracking performance on the (a) Simulation Rollouts and (b) Retargeted Motion Capture datasets. Each plot shows the resulting state trajectory (base height and pitch angle) after using a numerical inverse dynamics solver to track the original plan from each method. Trajectories are terminated and marked with an ’x’ upon failure (body height << 0.2 m or the vertical component of the base’s z-axis << 0.5). Most baselines produce untrackable plans that fail due to either the accumulation of physical errors (SAE) or inherent plan instability (high TRE). In contrast, only DynaFlow consistently generates trackable and stable trajectories.
@@ V-A1 Evaluation Metrics
To measure physical consistency, we use the Statewise Admissibility Error (SAE), introduced in DDAT [23]. For a given transition from $x_{i}$ to a ${x}_{i+1}$, the SAE is the Euclidean distance between ${x}_{i+1}$ and the closest reachable state under the system dynamics $f(x_{i},u)$ for any valid action $u\in\mathcal{U}$. The optimal action that achieves this closest state is obtained by solving the following inverse dynamics problem:
@@
To evaluate distributional similarity, we employ the Trajectory Reconstruction Error (TRE), which measures the mean squared error between a reference trajectory $\tilde{X}$ and a generated trajectory $\hat{X}$, which both start from the same initial state $x_{0}$.
$\text{TRE}(\hat{X},\tilde{X}):=\frac{1}{H-1}\sum_{i=1}^{H-1}\|\hat{x}_{i}-\tilde{x}_{i}\|_{2}^{2}.$
@@ V-A3 Results
On the dynamically feasible Simulation Rollouts dataset, the box plot shows that DynaFlow achieves a near-zero SAE, similar to SA-Rollout which benefits from ground-truth action labels. We note that this minimal non-zero SAE reflects the precision limit of the numerical inverse dynamics solver for evaluation rather than a model failure. In contrast, other state-predictive baselines like Vanilla State Flow and SA-State introduce noticeable feasibility errors, revealing that generative models can violate physical laws even when trained on ideal data. In terms of distributional similarity, DynaFlow attains a TRE comparable to the strongest baselines. The tracking experiment in Fig. 3 clearly demonstrates the consequence of these metrics: even the small physical errors from methods like Vanilla State Flow accumulate, causing the controller to fail. Conversely, the strict physical consistency (near-zero SAE) of DynaFlow’s trajectories enables stable and successful locomotion.
The performance gap becomes more pronounced on the physically inconsistent Retargeted Motion Capture dataset. This scenario invalidates action-supervised methods like State Action Flow. As seen in the box plot, this dataset has a high intrinsic SAE, which methods like Vanilla State Flow simply replicate. While corrective approaches (Guidance-based, Projection-based) offer partial improvements, they fail to fully resolve the infeasibility. The Tracker-based method successfully reduces the SAE to near-zero, but at the cost of a drastically increased TRE, indicating a significant deviation from the reference motion’s structure. In this context, DynaFlow is the only method that maintains both low SAE and low TRE simultaneously. The tracking results starkly reflect this. The high SAE of most baselines makes their plans untrackable from the start. Notably, even the Tracker-based method fails, as its physically plausible but structurally incoherent plan (high TRE) is too unstable to follow. Only DynaFlow, by satisfying both physical consistency and distributional similarity, produces a plan that can be successfully tracked, preserving the stylistic character of the motion.

### v3

@@ Abstract
This paper introduces DynaFlow, a novel framework that embeds a differentiable simulator directly into a flow matching model. By generating trajectories in the action space and mapping them to dynamically feasible state trajectories via the simulator, DynaFlow ensures all outputs are physically consistent by construction. This end-to-end differentiable architecture enables training on state-only demonstrations, allowing the model to simultaneously generate physically consistent state trajectories while inferring the underlying action sequences required to produce them. We demonstrate the effectiveness of our approach through quantitative evaluations and showcase its real-world applicability by deploying the generated actions onto a physical Go1 quadruped robot. The robot successfully reproduces diverse gaits present in the dataset, executes long-horizon motions in open-loop control and translates infeasible kinematic demonstrations into dynamically executable stylistic behaviors. These hardware experiments validate that DynaFlow produces deployable, highly effective motions on real-world hardware from state-only demonstrations, effectively bridging the gap between kinematic data and real-world execution.
@@ I Introduction
However, directly applying these generative models to character animation and robot control presents significant challenges. A primary limitation is the lack of physical consistency. Most generative models learn a statistical approximation of the data distribution from a finite set of examples, rather than the underlying physical principles governing the data. Consequently, there is no guarantee that the generated outputs will adhere to intrinsic physical principles or dynamic constraints. This means that when generating novel behaviors, the resulting motions can be either physically implausible, with artifacts like ground penetration, character floating, and foot sliding, or dynamically inconsistent, making them unsuitable to execute in the physical world. The issue becomes particularly pronounced when models are trained on datasets with inherent physical inconsistencies, such as those sourced from motion capture or generated via kinematic retargeting.
Another major hurdle is the scarcity of action data. While kinematic state trajectories are abundant, the corresponding hardware-specific action sequences (e.g., joint torques), necessary for deployment, are costly to collect. To bridge this gap, prior diffusion-based control approaches commonly employ a hierarchical framework, where the generative model produces desired state trajectories that are subsequently executed by a low-level tracking controller. Since dynamic feasibility is not enforced at the generative stage, stable execution hinges on additional adaptation between the planner and controller. Despite various efforts, mitigating this mismatch remains a significant challenge [12]. An alternative strategy follows a multi-expert distillation paradigm, where skill-specific expert controllers are first trained and their rollouts are distilled into a unified generative model. However, acquiring high-quality action experts is itself cumbersome, particularly when scaling to a diverse set of skills. Furthermore, the distillation stage introduces additional data engineering overhead to fully capture the robustness of these experts.
@@
To overcome these limitations, we propose DynaFlow, a novel dynamics-embedded flow matching framework that guarantees physically consistent motion generation while inferring deployable actions directly from state-only demonstrations. The core idea of DynaFlow is to integrate a differentiable simulator at the output of the flow matching prediction module. This simulator layer acts as a mapping from the space of action trajectories to the space of dynamically feasible state trajectories, ensuring that the model’s output strictly adheres to the laws of physics by construction. Furthermore, its differentiable nature enables end-to-end training via analytical gradients. During optimization, the model naturally discovers the action trajectory required to reconstruct a given state trajectory in a dynamically consistent manner. Consequently, DynaFlow alleviates the need for auxiliary tracking controllers or complex multi-stage distillation pipelines.
@@
We conduct experiments to validate the effectiveness of DynaFlow. Our quantitative analysis compares DynaFlow against several baselines, evaluating both dynamic feasibility and distributional similarity on two distinct datasets: a rich, strictly feasible dataset of quadruped locomotion and a challenging single-trajectory dataset from retargeted motion capture. Our results demonstrate that DynaFlow consistently generates strictly feasible trajectories, even when trained on a physically inconsistent dataset, while remaining competitive in distributional similarity. To showcase its real-world applicability, we deploy action trajectories generated by DynaFlow on a physical Go1 quadruped robot. The robot successfully reproduces diverse gaits observed in the training data and executes long-horizon motions with high accuracy in challenging open-loop experiments, validating the precision and coherence of the generated actions. Furthermore, we demonstrate its ability to translate infeasible retargeted motions into executable behaviors on hardware, bridging the gap between kinematic data and real-world execution.
@@
  • Dynamics-Embedded Generative Model: We propose DynaFlow, a generative model that embeds a differentiable simulator into a flow matching framework to enforce dynamically consistent trajectory generation.
  • Action-Free Learning via Analytical Gradients: Our method overcomes the scarcity of action data by leveraging analytical gradients to learn actions directly from state-only demonstrations.
  • Real-World Validation: We demonstrate the practical viability of our approach through hardware deployment on a physical quadruped robot, demonstrating successful real-world execution of the generated motions.
@@ II-A Generative Models for Control Tasks
The first paradigm is the hierarchical approach, where the generative model serves as a high-level state planner, while a separate low-level control module produces executable actions. This factorized design decouples planning and control, allowing the generative model to focus on state trajectory generation while other control modules handle deployment. A key design choice in this paradigm is how to implement the control module. For example, Decision Diffuser [1] uses ground-truth action labels to train an inverse dynamics network via supervised learning. In the absence of action labels, one common approach is to pair the planner with a reinforcement learning-based tracking controller [33, 31]. Alternatively, model-based approaches combine generative models with optimization-based controllers, where diffusion models either generate intermediate subgoals that are tracked by Model Predictive Control (MPC)[13] or provide initial guesses that are refined by trajectory optimization[25]. While effective, these approaches rely on auxiliary low-level tracking modules that are sensitive to out-of-distribution plans [12] or require extensive task-specific tuning, limiting scalability.
The second paradigm is the direct generation approach, where the generative model itself infers executable control inputs. Models in this category either model action trajectories directly or jointly model state–action pairs. This line of work has shown notable success across diverse domains, including manipulation [6, 10, 2], legged locomotion[11, 23], character control[12, 38, 37], and humanoid whole-body control[36]. While some works focus on generating action trajectories[6, 10, 2, 11, 37], others adopt joint formulations over state–action pairs [12] or couple states with latent action representations [38, 36]. Such formulations alleviate the challenge of action sequence conditioning and further demonstrate how task-level guidance can steer generated actions for downstream tasks. Despite these advances, the reliance on ground-truth action datasets poses a practical limitation, as such data is costly to obtain and often requires training or fine-tuning of expert controllers for each task[37].
@@ II-B Physics-Constrained Motion Generation
Our framework guarantees physical consistency by construction, embedding the simulator into the generative process, removing the need for external projection or guidance.
@@ II-C Differentiable Physics
The development of differentiable simulators[8, 7] has provided powerful tools for incorporating underlying physics directly into the learning process. By leveraging analytical gradients, these simulators enable sample-efficient training and provide interpretable learning signals from states. Building on this capability, several works have proposed policy learning methods that utilize gradient from differentiable simulators to guide policy optimization[22, 39]. Another line of research leverages the ability to backpropagate from states to actions, reformulating motion or skill imitation as a state-matching problem. DiffMimic[26] demonstrated efficient learning of dynamic character control from a single demonstration, while ILD[5] introduced a bidirectional state-matching objective that stabilizes training. In addition, differentiable simulation has been combined with state-diffusion planning to train latent action representations[27], which are then employed as tracking controllers for language-guided character control.
Our framework directly embeds a differentiable simulator into the flow matching framework, optimizing actions to match a target probability flow in the state space. This tight integration enables end-to-end learning of physically-grounded generative policy directly from state-only demonstrations.
@@ IV Dynamics-embedded Flow Matching
Our proposed framework, DynaFlow, guarantees the dynamic feasibility of the generated trajectory by embedding a simulator as a deterministic mapping module at the output, which projects preceding outputs into the space of physically valid state trajectories, ensuring feasibility by construction.
@@ IV-A Differentiable Physics as a Dynamics-Aware Embedding Layer
  • Action Prediction Network. A network $D_{\theta}$ predicts a sequence of future actions $U$ conditioned on the current noisy state trajectory $X_{t}$, conditioning $c$ including observations, command signals, and the time step $t$:
@@ IV-B Training Objective
Here, the expectation is over the time $t\sim\mathcal{U}(0,1)$, the base distribution $X_{0}\sim p_{0}$, and the target data distribution represented by the random variable $X_{1}\sim p_{1}$. $X_{t}$ denotes their linear interpolation, and $W$ is a weighting mask applied element-wise to balance contributions across state dimensions and timesteps. In practice, we approximate this expectation by sampling an expert trajectory $\tilde{X}_{1}\sim\mathcal{D}$ and generate a corresponding prediction $\hat{X}_{1}=\Phi(x_{0},D_{\theta}(X_{t},c,t))$.
@@ IV-D Datasets
We use two distinct datasets to evaluate performance under both controlled and challenging conditions. The first, Simulation Rollouts Dataset, comprises about 12,000 5-second trot and bound trajectories generated by rolling out a pre-trained policy [21] under various velocity commands. To establish a strictly feasible baseline for quantitative analysis, we re-generated the Simulation Rollouts Dataset in MuJoCo [35] under identical conditions, yielding dynamically feasible state–action pairs for controlled evaluation. The second, Retargeted Motion Capture Dataset, is a more challenging state-only dataset consisting of a single 2.54-second galloping trajectory. The motion was created by adapting German Shepherd motion capture data [41] to a Go1 robot skeleton, following the retargeting procedure of Peng et al. [24]. As kinematic retargeting does not incorporate dynamic properties or enforce actuation constraints, the resulting trajectory is often dynamically infeasible, exhibiting physical artifacts such as ground penetration. This makes it a demanding benchmark of model robustness and reflects real-world scenarios where only sparse kinematic data are available. The contrasting properties of these datasets are summarized in Table I.
TABLE I: Evaluation dataset characteristics.
Dataset Data Size Feasibility
Simulation Rollouts State+Action 12k traj. (16.7h) O
Retargeted MoCap State only 1 traj. (2.54s) X
@@ IV-E Implementation Details
Our framework is implemented in JAX[4], with MuJoCo XLA[7] serving as the differentiable simulation backend.
@@
We use a 1D Diffusion Transformer (DiT) backbone with three transformer blocks (10.3M parameters), conditioned with sinusoidal positional encodings over the horizon, MLP-based timestep embeddings, and task-specific attributes $c=[\,g,v,\omega,q,\dot{q},z_{\mathrm{gait}},v_{x}^{\mathrm{cmd}},v_{y}^{\mathrm{cmd}},\omega_{z}^{\mathrm{cmd}}]$. This vector includes gravity, base velocities, joint states, gait mode, and velocity commands. We train with Adam[15] optimizer with a learning rate of $2{\times}10^{-4}$, and apply Exponential Moving Averaging (EMA) with decay $0.995$ for 25k steps ($\approx$ 3h) on a single NVIDIA RTX 4090 GPU with a batch size of 1024.
@@ V-A Comparative Analysis on Physical Consistency and Distributional Similarity
Figure 3: Comparison of tracking performance on the (a) Simulation Rollouts and (b) Retargeted Motion Capture datasets. Each plot shows the resulting state trajectory (base height and pitch angle) after using a numerical inverse dynamics solver to track the original plan from each method. Trajectories are terminated and marked with an ‘x’ upon failure (body height << 0.2 m or the vertical component of the base’s z-axis << 0.5). Most baselines produce untrackable plans that fail due to either the accumulation of physical errors (SAE) or inherent plan instability (high TRE). In contrast, only DynaFlow consistently generates trackable and stable trajectories.
@@ V-A1 Evaluation Metrics
To measure physical consistency, we use the Statewise Admissibility Error (SAE), introduced in DDAT [3]. For a given transition from $x_{i}$ to ${x}_{i+1}$, the SAE is the Euclidean distance between ${x}_{i+1}$ and the closest reachable state under the system dynamics $f(x_{i},u)$ for any valid action $u\in\mathcal{U}$. The optimal action that achieves this closest state is obtained by solving the following inverse dynamics problem:
@@
To evaluate distributional similarity, we employ the Trajectory Reconstruction Error (TRE), which measures the mean squared error between a reference trajectory $\tilde{X}$ and a generated trajectory $\hat{X}$, with the same initial state $x_{0}$.
@@ V-A3 Results
On the dynamically feasible Simulation Rollouts dataset, the box plot shows that DynaFlow achieves near-zero SAE, similar to SA-Rollout, which benefits from ground-truth action labels. We note that this minimal non-zero SAE reflects the precision limit of the numerical inverse dynamics solver for evaluation rather than a model failure. In contrast, other state-predictive baselines like Vanilla State Flow and SA-State introduce noticeable feasibility errors, revealing that generative models can violate physical laws even when trained on ideal data. In terms of distributional similarity, DynaFlow attains a TRE comparable to the strongest baselines. The tracking experiment in Fig. 3 clearly demonstrates the consequence of these metrics: even small physical errors from methods like Vanilla State Flow accumulate, causing the controller to fail. Conversely, the strict physical consistency of DynaFlow’s trajectories enables stable locomotion.
The performance gap becomes more pronounced on the physically inconsistent Retargeted Motion Capture dataset. This scenario invalidates action-supervised methods like State Action Flow. As seen in the box plot, the dataset exhibits a high intrinsic SAE, which methods like Vanilla State Flow simply replicate. While corrective approaches (Guidance-based, Projection-based) offer partial improvements, they fail to fully resolve the infeasibility. The Tracker-based method successfully reduces the SAE to near-zero, but at the cost of a drastically increased TRE, indicating a significant deviation from the reference motion’s structure. In this context, DynaFlow is the only method that maintains both low SAE and low TRE simultaneously. The tracking results starkly reflect this. The high SAE of most baselines makes their plans untrackable from the start. Notably, even the Tracker-based method fails, as its physically plausible but structurally incoherent plan (high TRE) is too unstable to follow. Only DynaFlow, by satisfying both physical consistency and distributional similarity, produces a plan that can be successfully tracked, preserving motion style.